The sum of three natural numbers is 15. The second number is 1 more than the first number. The sum of the first and third numbers is equal to twice the second number. Find all three numbers.
Answer by richwmiller(17219)

4,5,6
x+1=y,
x+y+z=15,
x+z=2y